OpenStudy (anonymous):

simplify. x+2 over x-1 times 2x-2 over 5x+10

OpenStudy (anonymous):

try atleast

OpenStudy (anonymous):

2/5

OpenStudy (lalaly):

\[\huge{\frac{\cancel{x+2}}{\cancel{x-1}} \times \frac{2 \cancel{(x-1)}}{5 \cancel{(x+2)}}}\]\[=\huge{\frac{2}{5}}\]

OpenStudy (anonymous):

thankyou:)

OpenStudy (anonymous):

dont just put all ques from a quiz here try doing them

OpenStudy (anonymous):

if u encounter a problem ask us

OpenStudy (anonymous):

i have tried doing them . ive been sitting here for 3 hrs at least. honestly..

OpenStudy (anonymous):

so learn doing them

OpenStudy (anonymous):

tried that too...

OpenStudy (anonymous):

their are plenty of algebra tutorials online